[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Tue Oct 04, 2016

HSU Shift Log nhc.tsbadmin at noaa.gov
Wed Oct 5 03:43:54 UTC 2016


HSU Shift Log for Tue Oct 04 2016 23:43:47 GMT-0400 (EDT)

Shift: R

Forecaster: SR Stewart

Technical Issues Summary:

ATCF issues: None operationally. However, the TCM wind tool did not update  
in TAFB after advisory comp was closed out on Matthew.

N-AWIPS/AWIPS2 issues: None.

Other Techincal issues: None.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20161005/e5b95ea3/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list